OpenStudy (anonymous):

The ratio of the angle measures in a quadrilateral is 1 : 4 : 5 : 6. Find the measure of the largest angle.

OpenStudy (anonymous):

1:4:5:6 1+4+5+6=16 360/16=22.5 1*22.5=22.5 is the smallest angle. 4*22.5=90 is the next angle 5*22.5=112.5 is the third angle 6*22.5=135 is the largest angle Proof: 22.5+90+112.5+135=360 360=360

OpenStudy (anonymous):

I forgot how to do that one, but now I remember. Thanks :)